# Ecology, Spatial Structure, and Selection Pressure Induce Strong Signatures in Phylogenetic Structure

### v1

@@ Introduction
Effective use of phylogenetic metrics in scenarios where exact phylogenies are not available requires understanding of potential confounding effects from inaccuracies introduced by reconstruction. However, it is in precisely such scenarios where the new capabilities to characterize evolutionary dynamics could have the largest impact; large-scale systems can produce an intractable quantity of data, making phylometrics valuable as summary statistics of the evolutionary process (dolson2020interpreting). In particular, methods are needed to handle large-scale artificial life systems where complete, perfect visibility is not feasible and evolution operates according to implicit, contextually-dependent fitness dynamics (moreno2022exploring; kojima2023implementation). Thus, an additional goal of this paper is to quantify the magnitude and character of bias that reconstruction error introduces. In particular, we calculate phylometric metrics across range of reconstruction accuracy levels, and report the level of accuracy necessary to attain metric readings statistically indistinguishable from ground truth.
@@ Methods
Across all experiments, each treatment comprised 50 replicates.
@@
Previous work with hereditary stratigraphy used UPGMA distance-based reconstruction techniques (moreno2022hereditary). Large-scale reconstructions required for these experiments necessitated development of a more efficient technique that did not require all pairs (i.e., distance comparison. To accomplish this, we devised an agglomerative tree building algorithm that works by successively adding leaf organism annotations and percolating them down from the tree root along the tree path of internal nodes consistent with their fingerprint sequence, then affixing them where common ancestry ends. This new tree-building approach reduced compute time from multiple hours to around 5 minutes in most cases. Supplementary Listing LABEL:lst:build_tree_trie provides source code with full implementation details, see (moreno2024analysis) for a more detailed discussion.
To assess the efficacy of the new agglomerative tree-building approach, we calculated all reconstructed trees’ quartet distance to their respective reference. Quartet distance ranges from 0 (between identical trees) to 0.75 (between random trees), providing in this case a measure of reconstruction error. As expected, this measure of reconstruction error varied significantly with resolution for trees across all evolutionary regimes (Kruskal-Wallis tests; all ; Supplementary Table LABEL:tab:reconstruction-error-comparisons-between-resolutions). Reconstruction error also varied significantly with evolutionary regime for each reconstruction resolution level (Kruskal-Wallis tests; all ; Supplementary Table LABEL:tab:reconstruction-error-comparisons-between-regimes).
For 3% and 1% resolutions, mean reconstruction error was less than 0.01 in all cases and at 10% resolution mean reconstruction error was less than 0.05 in all cases. At 33% resolution, mean reconstruction error was less than 0.12 in all cases. The largest reconstruction errors observed at 1%, 3%, 10%, and 33% resolutions were, respectively, 0.051 (weak selection regime), 0.093 (weak 4 niche ecology regime), 0.14 (plain evolutionary regime), and 0.45 (plain evolutionary regime). Supplementary Table LABEL:tab:tree-reconstruction-quality-quartet-summary-statistics reports mean, median, standard deviation, and maxima for reconstruction error across surveyed evolutionary conditions.
@@
As a first step to characterizing the phylogenetic impact of spatial structure, ecology, and selection pressure, we tested whether surveyed evolutionary conditions exhibited detectable differences in a representative suite of four phylometrics: evolutionary distinctiveness, Colless-like index, mean pairwise distance, and sum pairwise distance. Figure 5 summarizes the distributions of each metric across surveyed conditions. Statistical tests confirmed that each phylometric exhibited significant variation among surveyed evolutionary conditions for both the simple model and Avida (Kruskal-Wallis tests; all ; per condition simple model, Avida; Supplementary Table LABEL:tab:phylostatistics-comparison-between-regimes-kwallis).
@@
We additionally performed a sensitivity analysis for results from the simple model over an alternate exponential mutation operator and earlier phylogeny sampling time points. The effects of evolutionary conditions on phylometrics were generally consistent with Figure 5(b) across surveyed conditions (Supplementary Figures LABEL:fig:perfect-tree-phylometrics-sensitivity-analysis and LABEL:fig:perfect-tree-phylometrics-heatmap-sensitivity-analysis).
@@
Figure 7 summarizes the distribution of surveyed phylometrics under the three surveyed ecological regimes and the control non-ecological regime, all with spatial population structure. Statistical tests confirmed that each phylometric exhibited significant variation among these evolutionary regimes, indicating the presence of detectable structural signatures in phylogenetic structure (Kruskal-Wallis tests; all ; per condition for simple model, Avida; Supplementary Tables LABEL:tab:phylostatistics-comparison-between-regimes-spatial-nuisance-kwallis and LABEL:tab:phylostatistics-comparison-between-regimes-spatial-nuisance-kwallis-avida).
@@
For these experiments, we again performed a sensitivity analysis over an alternate exponential mutation operator and earlier phylogeny sampling time points. We found the effects of evolutionary conditions on phylometrics to be generally consistent across surveyed conditions (Supplementary Figure LABEL:fig:perfect-tree-phylometrics-with-spatial-nuisance-sensitivity-analysis and Supplementary Tables LABEL:tab:phylostatistics-comparison-between-regimes-spatial-nuisance-kwallis and LABEL:tab:phylostatistics-comparison-between-resolutions-allpairs-wilcox-spatial-nuisance).
@@
Significant, large-effect changes were detected across all four phylometrics in each treatment in Gen3sis (see Figure 8). However, with the exception of sum pariwise distance, effect signs of treatments were opposite to those for individual-level phylogenies from Avida and the simple model across all phylometrics. Tip count effects seem likely to play a role in this discrepancy. Unlike Avida and the simple model, which held population size constant, species richness grew freely under the Gen3sis model. It is also possible that phylometric outcomes may be sensitive to granularity level of the taxonomic unit of the phylogeny. As noted in Section Methods and shown in Figure 9, in Avida experiments, phylogenies using genome-level tracking (as opposed to individual-level tracking) were notably different than those using individual-level tracking. These differences also included changes in the sign of some treatment effects, lending credence to the idea that differences between Gen3sis and the individual-level phylogenies could be partially the result of Gen3sis having a more abstract taxonomic unit.
@@
To explore this question, we compared phylometrics computed on reconstructed trees to corresponding true reference trees under the simple model (Wilcoxon tests; per condition; Supplementary Table LABEL:tab:phylostatistics-comparison-between-resolutions-allpairs-wilcox). To err towards conservatism in detecting phylometric biases, we did not correct for multiple comparisons. Reconstructions were performed across a range of precisions, ranging from 1% relative resolution for MRCA estimates (most precise) to 33% relative resolution for MRCA estimates (least precise). Precision was manipulated by adjusting the information content of underlying hereditary stratigraphic genome annotations used to perform phylogenetic reconstruction (moreno2022hereditary). Note that important differences exist the between nature of reconstruction error under hereditary stratigraphy versus traditional biosequence-based methods, discussed further below.
@@
Where detectable, estimation uncertainty bias decreased all surveyed phylometrics’ numerical value. So, when testing for expected increases in phylometric values, the potential for systematic false positives due to reconstruction error can be discounted. Supplementary Figure LABEL:fig:reconstructed-tree-phylometrics provides a full comparison of the distribution of phylometric estimates on reference trees with the distributions of phylometric estimates for reconstructed trees across reconstruction resolutions.
The relationship between reconstruction error and phylometric bias was similar under spatially structured regimes, the alternate exponential mutation operator, at earlier phylogeny sampling time points, and in the Avida model system (Supplementary Figures LABEL:fig:reconstructed-tree-phylometrics-error-sensitivity-analysis, LABEL:fig:reconstructed-tree-phylometrics-error-spatial-nuisance, LABEL:fig:reconstructed-tree-phylometrics-with-spatial-nuisance, LABEL:fig:reconstructed-tree-phylometrics-avida and LABEL:fig:reconstructed-tree-phylometrics-error-avida; Supplementary Table LABEL:tab:phylostatistics-comparison-between-resolutions-allpairs-wilcox and LABEL:tab:phylostatistics-comparison-between-resolutions-allpairs-wilcox-spatial-nuisance). Notably, however, reconstruction bias persisted at even 1% relative resolution in some conditions of the sensitivity analysis and Avida experiments. Forthcoming work has found that the byte-differentia hereditary stratigraphy configuration used for this experiment tends to lump closely contemporaneous lineage splitting events into a polytomy rather than a double-branching event — i.e., indicating uncertainty rather than introducing error. In contrast, reconstructions on single-bit differentia contain erroneously- sequenced branching rather than artifactual polytomies. So, future work should explore whether working with single bit differentia could lessen phylometric bias.
@@
Our last objective was to assess how reconstruction error might affect detection of phylometric signatures induced by treatment conditions. That is, we sought to perform a sort of “integration test” for detection of treatment conditions when working with imperfect reconstructions rather than perfect phylogenies. To this end, we compared the phylometric outcomes of strong/weak selection, spatial structure, ecology, and weak/rich ecology relative to plain conditions for phylogenies reconstructed at 1%, 3%, 10%, and 33% resolution levels. Supplemental Figures LABEL:fig:reconstructed-tree-phylometrics-progressive-heatmap and LABEL:fig:reconstructed-tree-phylometrics-progressive-heatmap-avida show heatmaps with sign, effect size, and significance of phylometric effects across gradations of reconstruction precision for the simple model and Avida, respectively. In most cases, 3% resolution sufficed to fully recover phylometric effects of treatments observed with perfectly-tracked phylogenies.

### v2

@@ Introduction
Effective use of phylogenetic metrics in scenarios calculated from estimated phylogenies (as opposed to exact phylogenies) requires consideration of potential confounding effects from inaccuracies introduced by reconstruction. However, it is in precisely such scenarios where the new capabilities to characterize evolutionary dynamics could have the greatest impact; large-scale systems can produce an intractable quantity of data, making phylometrics valuable as summary statistics of the evolutionary process (dolson2020interpreting). In particular, methods are needed to handle large-scale artificial life systems where complete, perfect visibility is not feasible and evolution operates according to implicit, contextually dependent fitness dynamics (moreno2022exploring; kojima2023implementation). Thus, an additional goal of this paper is to quantify the magnitude and character of bias introduced by reconstruction error from hereditary stratigraphy. To this end, we calculate phylometric metrics across a range of reconstruction accuracy levels, and report the level of accuracy necessary to attain metric readings statistically indistinguishable from ground truth.
This work builds on a series of recent studies developing, testing, and applying hereditary stratigraphy methodology since its introduction in moreno2022hereditary. We utilized the open-source hstrat software library’s hereditary stratigraphy algorithm implementations, which are publicly available via the Python Packaging Index (moreno2022hstrat). In perfect-tracking experiments, we collected ground-truth baseline phylogenies using the Phylotrack library (dolson2024phylotrack). moreno2024guide characterized trade-offs between memory use, inference precision, and inference accuracy across hereditary stratigraphy configurations, providing a foundation for best practices in applying the methodology. The present paper focuses on establishing foundations for using herediatry stratigraphy to infer evolutionary dynamics from phylogenetic history. Specifically, the work extends a conference paper moreno2023toward, by adding (1) replications of experiments in full-fledged evolution simulation frameworks (i.e., Avida and Gen3sis, introduced in methods) and (2) refining the set of phylogeny metrics employed. Findings from present work have facilitated application of hereditary stratigraphy to characterize dynamics in evolution simulations run on the 850,000-core Cerebras Wafer-Scale Engine (moreno2024trackable), which required engineering a simpler and more efficient algorithmic basis for hereditary stratigraphy (moreno2024structured). Although present work considers non-hybridizing phylogenies (i.e., asexual ancestry trees and species trees), methods applying hereditary stratigraphy to sexual populations have been proposed in moreno2024methods.
@@ Methods
Across all experiments, each treatment comprised 50 independent replicates. Testing over repeat simulations allowed us to characterize the influence of evolutionary dynamics on phylogenetic structure relative to background stochasticity, including the amount of distributional overlap in phylogeny metrics between treatments (i.e., Cliff’s delta statistic, discussed below).
@@
Previous work with hereditary stratigraphy used UPGMA distance-based reconstruction techniques (moreno2022hereditary). Large-scale reconstructions required for these experiments necessitated development of a more efficient technique that did not require all pairs (i.e., ) distance comparison. To accomplish this, we devised an agglomerative tree building algorithm that works by successively adding leaf organism annotations and percolating them down from the tree root along the tree path of internal nodes consistent with their fingerprint sequence, then affixing them where common ancestry ends. This new tree-building approach reduced compute time from multiple hours to around 5 minutes in most cases. Supplementary Listing LABEL:lst:build_tree_trie provides source code with full implementation details, see (moreno2024analysis) for a more detailed discussion.
To assess the efficacy of the new agglomerative tree-building approach, we calculated all reconstructed trees’ quartet distance to their respective reference. Quartet distance ranges from 0 (between identical trees) to 0.75 (between random trees), providing in this case a measure of reconstruction error. As expected, this measure of reconstruction error varied significantly with resolution for trees across all evolutionary regimes (Kruskal-Wallis tests; all ; Supplementary Table S1). Reconstruction error also varied significantly with evolutionary regime for each reconstruction resolution level (Kruskal-Wallis tests; all ; Supplementary Table S1).
For 3% and 1% resolutions, mean reconstruction error was less than 0.01 in all cases and at 10% resolution mean reconstruction error was less than 0.05 in all cases. At 33% resolution, mean reconstruction error was less than 0.12 in all cases. The largest reconstruction errors observed at 1%, 3%, 10%, and 33% resolutions were, respectively, 0.051 (weak selection regime), 0.093 (weak 4 niche ecology regime), 0.14 (plain evolutionary regime), and 0.45 (plain evolutionary regime). Supplementary Table 27 reports mean, median, standard deviation, and maxima for reconstruction error across surveyed evolutionary conditions.
@@
As a first step to characterizing the phylogenetic impact of spatial structure, ecology, and selection pressure, we tested whether surveyed evolutionary conditions exhibited detectable differences in a representative suite of four phylometrics: evolutionary distinctiveness, Colless-like index, mean pairwise distance, and sum pairwise distance. Figure 5 summarizes the distributions of each metric across surveyed conditions. Statistical tests confirmed that each phylometric exhibited significant variation among surveyed evolutionary conditions for both the simple model and Avida (Kruskal-Wallis tests; all ; per condition simple model, Avida; Supplementary Table 27).
@@
We additionally performed a sensitivity analysis for results from the simple model over an alternate exponential mutation operator and earlier phylogeny sampling time points. The effects of evolutionary conditions on phylometrics were generally consistent with Figure 5(b) across surveyed conditions (Supplementary Figures 12 and 20).
@@
Figure 7 summarizes the distribution of surveyed phylometrics under the three surveyed ecological regimes and the control non-ecological regime, all with spatial population structure. Statistical tests confirmed that each phylometric exhibited significant variation among these evolutionary regimes, indicating the presence of detectable structural signatures in phylogenetic structure (Kruskal-Wallis tests; all ; per condition for simple model, Avida; Supplementary Tables 27 and 27).
@@
For these experiments, we again performed a sensitivity analysis over an alternate exponential mutation operator and earlier phylogeny sampling time points. We found the effects of evolutionary conditions on phylometrics to be generally consistent across surveyed conditions (Supplementary Figure 19 and Supplementary Tables 27 and 27).
@@
Significant, large-effect changes were detected across all four phylometrics in each treatment in Gen3sis (see Figure 8). However, with the exception of sum pariwise distance, effect signs of treatments were opposite to those for individual-level phylogenies from Avida and the simple model across all phylometrics. Tip count effects seem likely to play a role in this discrepancy. Unlike Avida and the simple model, which held population size constant, species richness grew freely under the Gen3sis model. It is also possible that phylometric outcomes may be sensitive to granularity level of the taxonomic unit of the phylogeny. As noted in our methods and shown in Figure 9, in Avida experiments, phylogenies using genome-level tracking (as opposed to individual-level tracking) were notably different than those using individual-level tracking. These differences also included changes in the sign of some treatment effects, lending credence to the idea that differences between Gen3sis and the individual-level phylogenies could be partially the result of Gen3sis having a more abstract taxonomic unit.
@@
To explore this question, we compared phylometrics computed on reconstructed trees to corresponding true reference trees under the simple model (Wilcoxon tests; per condition; Supplementary Table 27). To err towards conservatism in detecting phylometric biases, we did not correct for multiple comparisons. Reconstructions were performed across a range of precisions, ranging from 1% relative resolution for MRCA estimates (most precise) to 33% relative resolution for MRCA estimates (least precise). Precision was manipulated by adjusting the information content of underlying hereditary stratigraphic genome annotations used to perform phylogenetic reconstruction (moreno2022hereditary). Note that important differences exist the between nature of reconstruction error under hereditary stratigraphy versus traditional biosequence-based methods, discussed further below.
@@
Where detectable, estimation uncertainty bias decreased all surveyed phylometrics’ numerical value. So, when testing for expected increases in phylometric values, the potential for systematic false positives due to reconstruction error can be discounted. Supplementary Figure 23 provides a full comparison of the distribution of phylometric estimates on reference trees with the distributions of phylometric estimates for reconstructed trees across reconstruction resolutions.
The relationship between reconstruction error and phylometric bias was similar under spatially structured regimes, the alternate exponential mutation operator, at earlier phylogeny sampling time points, and in the Avida model system (Supplementary Figures 21, 25, 24, 26 and 27; Supplementary Table 27 and 27). Notably, however, reconstruction bias persisted at even 1% relative resolution in some conditions of the sensitivity analysis and Avida experiments. Forthcoming work has found that the byte-differentia hereditary stratigraphy configuration used for this experiment tends to lump closely contemporaneous lineage splitting events into a polytomy rather than a double-branching event — i.e., indicating uncertainty rather than introducing error. In contrast, reconstructions on single-bit differentia contain erroneously- sequenced branching rather than artifactual polytomies. So, future work should explore whether working with single bit differentia could lessen phylometric bias.
@@
Our last objective was to assess how reconstruction error might affect detection of phylometric signatures induced by treatment conditions. That is, we sought to perform a sort of “integration test” for detection of treatment conditions when working with imperfect reconstructions rather than perfect phylogenies. To this end, we compared the phylometric outcomes of strong/weak selection, spatial structure, ecology, and weak/rich ecology relative to plain conditions for phylogenies reconstructed at 1%, 3%, 10%, and 33% resolution levels. Supplemental Figures 30 and 31 show heatmaps with sign, effect size, and significance of phylometric effects across gradations of reconstruction precision for the simple model and Avida, respectively. In most cases, 3% resolution sufficed to fully recover phylometric effects of treatments observed with perfectly-tracked phylogenies.
@@ Acknowledgment
S1 Supplemental Material
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head= Tree Distance Metric Statistic Reconstruction Fidelity Epoch Mutation Distribution
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
head to column names,
respect all,
[…]
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head= Tree Distance Metric Statistic Evolutionary Regime Epoch Mutation Distribution
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
head to column names,
respect all,
[…]
Generations Ago (approx.)
Generations Ago (approx.)
200,000
150,000
100,000
50,000
0
200,000
150,000
100,000
50,000
0
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head=evolutionary regime reconstruction quality sensitivity analysis conditions quartet dist mean quartet dist median quartet dist std quartet dist max
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
head to column names,
respect all,
[…]
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head=evolutionary regime reconstruction quality sensitivity analysis conditions triplet dist mean triplet dist median triplet dist std triplet dist max
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
head to column names,
respect all,
[…]
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head=evolutionary regime 1 evolutionary regime 2 Phylometric Statistic Epoch Mutation Distribution
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
head to column names,
respect all,
[…]
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head=evolutionary regime 1 evolutionary regime 2 Phylometric Statistic Epoch Mutation Distribution
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
head to column names,
respect all,
[…]
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head=Reconstruction Fidelity Phylometric Statistic Evolutionary Regime Epoch Mutation Distribution
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
head to column names,
respect all,
[…]
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head=Reconstruction Fidelity Phylometric Statistic Evolutionary Regime Epoch Mutation Distribution
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
head to column names,
respect all,
[…]
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head= Phylometric Statistic Epoch Mutation Distribution
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
after reading=
head to column names,
respect all,
[…]
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head= Phylometric Statistic Epoch Mutation Distribution
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
after reading=
head to column names,
respect all,
[…]
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head= Phylometric Statistic Epoch Mutation Distribution
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
head to column names,
respect all,
[…]
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head= Phylometric Statistic Epoch Mutation Distribution
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
head to column names,
respect all,
[…]
S1.1 Trie-based Phylogenetic Tree Reconstruction Implementation
This section provides source code and docstrings for trie-based phylogeny reconstruction.
The main reconstruction logic takes place in build_trie_from_artifacts (Listing LABEL:lst:build_tree_from_artifacts). First, hereditary stratigraphic annotations associated with each extant population member are sorted by the generational depth of their lineages in ascending order. This ensures that higher-per-generation-density hereditary stratigraphic records of annotations from shorter lineages are encountered first. (In this work, because generations are synchronous, this step had no effect.)
Next, annotations are inserted into a trie data structure one at a time. All annotations begin at the root node of the trie, representing a universal common ancestor. Each subsequent node in the trie denotes a particular “fingerprint” value associated with its generation of creation. (This is referred to in the source code as an “allele”). Paths traced from the root node towards the tips therefore represent a particular sequence of allele generation events; organisms will share common allele generation histories for the over the evolutionary interval they share common ancestry.
Each annotation descends the trie along the path corresponding to the sequence of alleles contained within its own record. The GetDeepestCongruousAllele member function of the TrieInnerNode class implements this process. When no further consistent alleles are present in the trie, the remaining alleles within the annotation are unrolled to create a new unifurcating branch. The label corresponding to the annotation is then appended as the leaf node on that branch. The InsertTaxon member function of the TrieInnerNode class implements this process. The TrieInnerNode and TrieLeafNode classes are provided in Listings LABEL:lst:TrieInnerNode and LABEL:lst:TrieLeafNode.
The build_tree_trie method (Listing LABEL:lst:build_tree_trie) serves as the API entrypoint for the overall tree reconstruction routine. Default options were used for all reconstructions in this work. The build_tree_trie entrypoint delegates to build_tree_trie_ensemble (Listing LABEL:lst:build_tree_trie_ensemble), which ultimately delegates to build_build_trie_from_artifacts. These additional source files are included due to docstring parameter descriptions referenced in other files and information about minor trie postprocessing performed to perform ancestor origin time estimates and convert the trie structure into a standard phylogenetic output format. Listing LABEL:lst:AssignOriginTimeNaiveTriePostprocessor provides the particular postprocessing implementation used to estimate ancestor origin times in reconstructions performed as part of this work.
The full source context for the files included in listings can be found at https://github.com/mmore500/hstrat.